# Joint Constraints on Fuzzy and Warm Dark Matter from Satellite Populations of the Milky Way and Andromeda

### v1

@@ 4 Modeling Framework
In this section, we introduce our modeling framework based on the galaxy–halo connection model (2020ApJ...893...48N), including the adopted dark matter-only zoom-in simulations, model parameters, and statistical inference procedure. More detailed description of the galaxy–halo connection model can be found in 2019ApJ...873...34N and 2020ApJ...893...48N.
4.1 Overview of the modeling framework
@@
Our goal is to constrain dark matter models using the observed populations of satellite galaxies. To achieve this, we first construct a model capable of predicting satellite populations to fit the observational data. In this work, we use high-resolution dark matter-only zoom-in simulations based on CDM to generate subhalo populations. We then apply the galaxy–halo connection model to populate these subhalos with satellite galaxies. The impact of different dark matter models is introduced through the suppression of the subhalo mass function in non-CDM model relative to CDM. By varying the model parameters, we can thus generate predicted satellite populations.
@@
It is important to note that our approach assumes that the radial distribution of satellites remains unchanged across different dark matter models. This assumption is justified for WDM (2014MNRAS.439..300L; 2017MNRAS.464.4520B), while it is more uncertain for FDM. However, since the mean mass of the faintest satellites used in our analysis exceeds $10^{8}~M_{\odot}$ (2020ApJ...893...48N), the wave interference effects of FDM can be safely neglected within the mass range of interest (2023MNRAS.524.4256M; 2025ApJ...986..127N). Therefore, we consider this assumption is also reasonable for FDM.
4.2 Zoom-in simulations
@@
We use high-resolution dark matter-only zoom-in simulations from 2015ApJ...810...21M, which are the same as those adopted in 2020ApJ...893...48N, to generate subhalo populations. The simulation suite contains 45 MW-mass host halos with virial masses between $1.18$ and $1.57\times 10^{12}~M_{\odot}$. Subhalos in these simulations are well resolved down to a present-day maximum circular velocity of $V_{\max}\approx 9~\mathrm{km\,s^{-1}}$. In addition, we require each subhalo to have a peak maximum circular velocity $V_{\mathrm{peak}}>10~\mathrm{km\,s^{-1}}$.
@@
Because the detailed merger history of the MW, particularly the early accretion of an galaxy with the Large Magellanic Cloud (LMC) mass, can affect its satellite population (2020MNRAS.495..743B), we follow 2020ApJ...893...48N and adopt two host halos with assembly histories similar to that of the MW, which have virial masses of $1.26$ and $1.57\times 10^{12}~M_{\odot}$, respectively.
@@
To avoid potential systematic uncertainties introduced by using different simulations, we use the same high-resolution dark matter-only zoom-in simulations from 2015ApJ...810...21M for M31, even though its virial mass and merger history remain highly uncertain (2018ApJ...864...34S; 2019ApJ...872...24V; 2023ApJ...948..104P; 2024MNRAS.528.2653Z; 2025arXiv250206948D). Specifically, we select two host halos with virial masses of $1.19$ and $1.35\times 10^{12}~M_{\odot}$ to account for the uncertainty in M31 assembly history. We have verified that our results are insensitive to this choice, as using different simulated host halos yields similar constraints. We discuss our treatment of the large uncertainty in M31 virial mass in Section 4.3.
4.3 Model parameters and statistical inference
@@
The galaxy–halo connection model includes eight parameters (2020ApJ...893...48N). The first two parameters describe the mapping between the subhalo peak maximum circular velocity $V_{\mathrm{peak}}$ and the satellite absolute magnitude $M_{V}$: the power-law slope of the luminosity function, $\alpha$, and the luminosity scatter, $\sigma_{M}$. The next two parameters characterize the galaxy formation efficiency using the subhalo peak virial mass $M$: the 50% galaxy occupation peak virial mass, $\mathcal{M}_{50}$, and the scatter in the galaxy occupation fraction, $\sigma_{\mathrm{gal}}$. The fifth parameter, $\mathcal{B}$, quantifies the strength of subhalo disruption due to baryonic effects. The remaining three parameters describe the mapping between the subhalo virial radius at accretion and the satellite half-light radius $r_{1/2}$: the amplitude of the galaxy–halo size relation, $\mathcal{A}$, the scatter in half-light radius at fixed halo size, $\sigma_{\log_{10}R}$, and the power-law index of the size relation, $n$. We also introduce the parameter $M_{\mathrm{hm}}$ to model the effects of FDM or WDM (2021PhRvL.126i1101N). Its impact enters through the subhalo mass function suppression described by Equations (8) or (12), and $M_{\mathrm{hm}}$ can be converted into the corresponding dark matter particle mass using Equations (7) or (11).
@@
Subhalo abundance is known to scale linearly with host halo mass (2012MNRAS.424.2715W; 2015ApJ...810...21M; 2025arXiv250912313T). In addressing the uncertainty in host halo mass, 2020ApJ...893...48N simply scale the final results based on the average mass of the two realistic MW-like simulations and the $2\,\sigma$ upper limit of the MW mass predicted in 2019MNRAS.484.5453C, yielding a very conservative result. In our work, we introduce the mass parameter $M_{\mathrm{host}}$, which represents the virial mass of the MW or M31. This parameter linearly scales $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ relative to the simulated host halo mass $M_{\mathrm{sim}}$, such that $\mathcal{M}_{50}=\mathcal{M}_{50}^{\mathrm{(sim)}}(M_{\mathrm{host}}/M_{\mathrm{sim}})$ and $M_{\mathrm{hm}}=M_{\mathrm{hm}}^{\mathrm{(sim)}}(M_{\mathrm{host}}/M_{\mathrm{sim}})$. We assign a log-normal prior within a $2\,\sigma$ range to the $M_{\mathrm{host}}$ parameter, so its role is to linearly scale $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ using the mass prior information, effectively accounting for the uncertainties in the MW and M31 masses.
@@
For the MW, we use the results from 2019MNRAS.484.5453C, corresponding to $\log_{10}(M_{\mathrm{MW}})$ in the range of [12.00, 12.26]. For M31, we adopt the baseline prior from 2024MNRAS.528.2653Z, corresponding to $\log_{10}(M_{\mathrm{M31}})$ in the range of [11.81, 12.45]. At the same time, we note that M31’s mass uncertainty is significant. The result from 2023ApJ...948..104P ($\sim 3.02\times 10^{12}~M_{\odot}$) is about a factor of 2-3 higher than our baseline setting from 2024MNRAS.528.2653Z ($\sim 1.33\times 10^{12}~M_{\odot}$). Therefore, we also consider the result from 2023ApJ...948..104P, corresponding to $\log_{10}(M_{\mathrm{M31}})$ in the range of [12.22, 12.74]. The priors for the other parameters are consistent with those in 2025ApJ...986..127N.
@@
To account for uncertainty in the observer’s location, we fix the observer’s distance to $8~\mathrm{kpc}$ for the MW (2020ApJ...893...48N) and to $776~\mathrm{kpc}$ for M31 (2022ApJ...938..101S), and then draw a random viewing position on the sphere at the corresponding radius. For MW realizations, we additionally rotate each simulated host halo, so that the LMC analog lies at the observed sky position of the LMC. For M31 realizations, we rotate the simulated system, so that the host center projects to the observed sky position of M31.
@@
Therefore, for the MW or M31, given these 10 parameters $\boldsymbol{\theta}$ and simulation data and using the binning method from 2025ApJ...986..127N, the number of predicted observable satellites in luminosity and surface brightness bin $i$ in a given realization is
@@
The probability of observing $N_{\mathrm{obs},i}$ satellites in bin $i$ is (2019ApJ...873...34N; 2020ApJ...893...48N)
@@
Before constructing the total likelihood function, we clarify that when jointly modeling the MW and M31 satellite populations, we treat $\mathcal{M}_{50}$, $\sigma_{\mathrm{gal}}$, and $M_{\mathrm{hm}}$ as shared parameters, while all other parameters are allowed to vary independently for the MW and M31. We make this choice because the mass of M31 is highly uncertain and may exceed the mass range covered by our simulation suite (2015ApJ...810...21M; 2023ApJ...948..104P; 2024MNRAS.528.2653Z). Although our galaxy–halo connection model is sufficiently flexible to reproduce the observed data even when the host mass lies outside the simulated range, the inferred parameters in such cases would be systematically biased (see Appendix A for a detailed discussion). In contrast, $\mathcal{M}_{50}$, $\sigma_{\mathrm{gal}}$, and $M_{\mathrm{hm}}$ can be reliably scaled according to the mass priors, yielding physically consistent values.
@@
We use emcee (2013PASP..125..306F) in the constraint process for the model parameters. Specifically, we run 112 walkers for $10^{5}$ steps, and retain more than $10^{4}$ independent samples overall to sample the posterior or probability distribution funtions (PDFs) of the parameters.
Figure 3: The PDF and contour maps (68% and 95% CLs) of the shared parameters in the FDM model. The shaded regions in the one-dimensional PDFs and the quoted parameter values correspond to 68% CL. The parameters $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ are in units of $M_{\odot}$, and $\sigma_{\mathrm{gal}}$ is in dex.
@@
5 Constraint results
5.1 Fuzzy dark matter
@@
In Figure 3, we show the marginalized PDFs and contour maps for the shared parameters in the FDM model. From the PDFs, we derive $M_{\mathrm{hm}}<3.10\times 10^{7}~M_{\odot}$ at 95% confidence level (CL), which can be converted into a lower limit on the FDM particle mass using Equation (7), yielding
$m_{\mathrm{FDM}}>1.75\times 10^{-20}~\mathrm{eV}\ (95\%\ \mathrm{CL}).$ (18)
@@
This constraint improves upon the result obtained using only the MW satellite population by 2025ApJ...986..127N, $m_{\mathrm{FDM}}>1.4\times 10^{-20}~\mathrm{eV}$, by about 25%. The precision of our limit is also comparable to that from the Lyman-$\alpha$ forest analysis in 2021PhRvL.126g1302R, which yields $m_{\mathrm{FDM}}>2.0\times 10^{-20}~\mathrm{eV}$.
@@
Besides, to obtain a more prior-independent constraint, we also adopt a posterior ratio of 20:1, corresponding to the point where the posterior probability drops to one-twentieth of its maximum value (2023MNRAS.524L..84P; 2024MNRAS.535.1652K; 2025PhRvD.111f3079T). Our lower prior bound on $M_{\mathrm{hm}}$, $10^{7}~M_{\odot}$, coincides with the region where the posterior distribution flattens out, and we have verified the robustness of the 20:1 posterior ratio result by varying the prior range. For the FDM case, this criterion yields $M_{\mathrm{hm}}<4.19\times 10^{7}~M_{\odot}$, corresponding to
$m_{\mathrm{FDM}}>1.41\times 10^{-20}~\mathrm{eV}\ (\text{20:1}\ \mathrm{posterior\ ratio}).$ (19)
@@
Our result is about three times stronger than the posterior odds ratio of 20:1 constraint derived from strong gravitational lensing by 2023MNRAS.524L..84P, who obtained $m_{\mathrm{FDM}}>4.4\times 10^{-21}~\mathrm{eV}$. We note that 2023MNRAS.524L..84P and 2025PhRvD.111f3079T adopt a posterior odds ratio, while 2021ApJ...917....7N use a likelihood ratio. However, because our posterior flattens toward the lower prior boundary and we adopt a uniform prior in $\log_{10}(M_{\mathrm{hm}})$, these criteria reduce to the same effective threshold as our posterior ratio.
@@
We present the full PDFs and contour maps of the MW and M31 parameters for the FDM case in Appendix A. We have verified that, aside from the shared parameters, the posterior distributions of the other parameters are consistent with those obtained from the separate MW and M31 fits. Furthermore, we rerun our model using the larger M31 mass prior from 2023ApJ...948..104P, and obtain similar and slightly weaker constraint results with $m_{\mathrm{FDM}}>1.69\times 10^{-20}\,\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.37\times 10^{-20}\,\mathrm{eV}$ (20:1 posterior ratio).
5.2 Warm dark matter

### v2

@@ IV Modeling Framework
In this section, we introduce our modeling framework based on the empirical galaxy–halo connection model (E. O. Nadler et al., 2020), including the adopted dark matter-only zoom-in simulations, model parameters, and statistical inference procedure. [^3] https://github.com/eonadler/subhalo_satellite_connection More detailed description of the empirical galaxy–halo connection model can be found in E. O. Nadler et al. (2019b) and E. O. Nadler et al. (2020).
IV.1 Overview of the modeling framework
@@
Our goal is to constrain dark matter models using the observed populations of satellite galaxies. To achieve this, we first construct a model capable of predicting satellite populations to fit the observational data. In this work, we use high-resolution dark matter-only zoom-in simulations based on CDM to generate subhalo populations. We then apply the empirical galaxy–halo connection model to populate these subhalos with satellite galaxies. The impact of different dark matter models is introduced through the suppression of the subhalo mass function in non-CDM model relative to CDM. By varying the model parameters, we can thus generate predicted satellite populations.
@@
It is important to note that our approach assumes that the radial distribution of satellites remains unchanged across different dark matter models. This assumption is justified for WDM (M. R. Lovell et al., 2014; S. Bose et al., 2017), while it is more uncertain for FDM. However, since the mean mass of the faintest satellites used in our analysis exceeds $10^{8}~M_{\odot}$ (E. O. Nadler et al., 2020), the wave interference effects of FDM can be safely neglected within the mass range of interest (S. May & V. Springel, 2023; E. O. Nadler et al., 2025). Therefore, we consider this assumption is also reasonable for FDM.
IV.2 Zoom-in simulations
@@
We use high-resolution dark matter-only zoom-in simulations from Y.-Y. Mao et al. (2015), which are the same as those adopted in E. O. Nadler et al. (2020), to generate subhalo populations. The simulation suite contains 45 MW-mass host halos with virial masses between $1.18$ and $1.57\times 10^{12}~M_{\odot}$. Subhalos in these simulations are well resolved down to a present-day maximum circular velocity of $V_{\max}\approx 9~\mathrm{km\,s^{-1}}$. In addition, we require each subhalo to have a peak maximum circular velocity $V_{\mathrm{peak}}>10~\mathrm{km\,s^{-1}}$.
@@
Because the detailed merger history of the MW, particularly the early accretion of a galaxy with the Large Magellanic Cloud (LMC) mass, can affect its satellite population (S. Bose et al., 2020), we follow E. O. Nadler et al. (2020) and adopt two host halos with assembly histories similar to that of the MW, which have virial masses of $1.26$ and $1.57\times 10^{12}~M_{\odot}$, respectively.
@@
To avoid potential systematic uncertainties introduced by using different simulations, we use the same high-resolution dark matter-only zoom-in simulations from Y.-Y. Mao et al. (2015) for M31, even though its virial mass and merger history remain highly uncertain (M. Semczuk et al., 2018; R. P. van der Marel et al., 2019; E. Patel & K. S. Mandel, 2023; X. Zhang et al., 2024; A. Doliva-Dolinsky et al., 2025). Specifically, we select two host halos with virial masses of $1.19$ and $1.35\times 10^{12}~M_{\odot}$ to account for the uncertainty in M31 assembly history. We have verified that our results are insensitive to this choice, as using different simulated host halos yields similar constraints. We discuss our treatment of the large uncertainty in M31 virial mass in Section IV.3.
IV.3 Model parameters and statistical inference
@@
The empirical galaxy–halo connection model includes eight parameters (E. O. Nadler et al., 2020). The first two parameters describe the mapping between the subhalo peak maximum circular velocity $V_{\mathrm{peak}}$ and the satellite absolute magnitude $M_{V}$: the power-law slope of the luminosity function, $\alpha$, and the luminosity scatter, $\sigma_{M}$. The next two parameters characterize the galaxy formation efficiency based on the subhalo peak virial mass $M$: the peak virial mass at which the galaxy occupation fraction is 50%, denoted as $\mathcal{M}_{50}$, and the parameter that determines the tilt of the galaxy occupation fraction, $\mathcal{S}_{\mathrm{gal}}$. The fifth parameter, $\mathcal{B}$, quantifies the strength of subhalo disruption due to baryonic effects. The remaining three parameters describe the mapping between the subhalo virial radius at accretion and the satellite half-light radius $r_{1/2}$: the amplitude of the galaxy–halo size relation, $\mathcal{A}$, the scatter in half-light radius at fixed halo size, $\sigma_{\log_{10}R}$, and the power-law index of the size relation, $n$. We also introduce the parameter $M_{\mathrm{hm}}$ to model the effects of FDM or WDM (E. O. Nadler et al., 2021b). Its impact enters through the subhalo mass function suppression described by Equations (8) or (12), and $M_{\mathrm{hm}}$ can be converted into the corresponding dark matter particle mass using Equations (7) or (11).
@@
Subhalo abundance is known to scale linearly with host halo mass (J. Wang et al., 2012; Y.-Y. Mao et al., 2015; C. Y. Tan et al., 2025b). In addressing the uncertainty in host halo mass, E. O. Nadler et al. (2020) simply scale the final results based on the average mass of the two realistic MW-like simulations and the $2\,\sigma$ upper limit of the MW mass predicted in T. M. Callingham et al. (2019), yielding a very conservative result. In our work, we introduce the mass parameter $M_{\mathrm{host}}$, which represents the virial mass of the MW or M31. This parameter linearly scales $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ relative to the simulated host halo mass $M_{\mathrm{sim}}$, such that $\mathcal{M}_{50}=\mathcal{M}_{50}^{\mathrm{(sim)}}(M_{\mathrm{host}}/M_{\mathrm{sim}})$ and $M_{\mathrm{hm}}=M_{\mathrm{hm}}^{\mathrm{(sim)}}(M_{\mathrm{host}}/M_{\mathrm{sim}})$. We assign a log-normal prior within a $2\,\sigma$ range to the $M_{\mathrm{host}}$ parameter, so its role is to linearly scale $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ using the mass prior information, effectively accounting for the uncertainties in the MW and M31 masses.
@@
For the MW, we use the results from T. M. Callingham et al. (2019), corresponding to $\log_{10}(M_{\mathrm{MW}})$ in the range of [12.00, 12.26]. For M31, we adopt the baseline prior from X. Zhang et al. (2024), corresponding to $\log_{10}(M_{\mathrm{M31}})$ in the range of [11.81, 12.45]. At the same time, we note that M31’s mass uncertainty is significant. The result from E. Patel & K. S. Mandel (2023) ($\sim 3.02\times 10^{12}~M_{\odot}$) is about a factor of 2-3 higher than our baseline setting from X. Zhang et al. (2024) ($\sim 1.33\times 10^{12}~M_{\odot}$). Therefore, we also consider the result from E. Patel & K. S. Mandel (2023), corresponding to $\log_{10}(M_{\mathrm{M31}})$ in the range of [12.22, 12.74]. Note that while the mass definitions in T. M. Callingham et al. (2019) and X. Zhang et al. (2024) differ from the one adopted in this work, we have consistently converted these values to ensure a uniform mass definition across our analysis. The priors for the other parameters are consistent with those in E. O. Nadler et al. (2025).
@@
To account for uncertainty in the observer’s location, we fix the observer’s distance to $8~\mathrm{kpc}$ for the MW (E. O. Nadler et al., 2020) and to $776~\mathrm{kpc}$ for M31 (A. Savino et al., 2022), and then draw a random viewing position on the sphere at the corresponding radius. For MW realizations, we additionally rotate each simulated host halo, so that the LMC analog lies at the observed sky position of the LMC. For M31 realizations, we rotate the simulated system, so that the host center projects to the observed sky position of M31.
@@
Therefore, for the MW or M31, given these 10 parameters $\boldsymbol{\theta}$ and simulation data and using the binning method from E. O. Nadler et al. (2025), the number of predicted observable satellites in luminosity and surface brightness bin $i$ in a given realization is
@@
The probability of observing $N_{\mathrm{obs},i}$ satellites in bin $i$ is (E. O. Nadler et al., 2019b, 2020)
@@
Before constructing the total likelihood function, we clarify that when jointly modeling the MW and M31 satellite populations, we treat $\mathcal{M}_{50}$, $\mathcal{S}_{\mathrm{gal}}$, and $M_{\mathrm{hm}}$ as shared parameters, while all other parameters are allowed to vary independently for the MW and M31. We make this choice because the mass of M31 is highly uncertain and may exceed the mass range covered by our simulation suite (Y.-Y. Mao et al., 2015; E. Patel & K. S. Mandel, 2023; X. Zhang et al., 2024). Although our galaxy–halo connection model is sufficiently flexible to reproduce the observed data even when the host mass lies outside the simulated range, the inferred parameters in such cases would be systematically biased (see Appendix A for a detailed discussion). In contrast, $\mathcal{M}_{50}$, $\mathcal{S}_{\mathrm{gal}}$, and $M_{\mathrm{hm}}$ can be reliably scaled according to the mass priors, yielding physically consistent values.
@@
We use emcee (D. Foreman-Mackey et al., 2013) in the constraint process for the model parameters. Specifically, we run 112 walkers with $10^{5}$ steps for each walker, and discard the first 10 maximum autocorrelation lengths as burn-in ( $\sim 10^{4}$ steps for each walker) to sample the posterior distributions or the probability distribution functions (PDFs) of the parameters.
Figure 3: The PDF and contour maps (68% and 95% CLs) of the shared parameters in the FDM model. The shaded regions in the one-dimensional PDFs and the quoted parameter values correspond to 68% CL. The parameters $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ are in units of $M_{\odot}$, and $\mathcal{S}_{\mathrm{gal}}$ is in dex.
@@
V Constraint results
V.1 Fuzzy dark matter
@@
In Figure 3, we show the marginalized PDFs and contour maps for the shared parameters in the FDM model. From the PDFs, we derive $M_{\mathrm{hm}}<3.12\times 10^{7}~M_{\odot}$ at 95% confidence level (CL), which can be converted into a lower limit on the FDM particle mass using Equation (7), yielding
$m_{\mathrm{FDM}}>1.74\times 10^{-20}~\mathrm{eV}\ (95\%\ \mathrm{CL}).$ (18)
@@
This constraint improves upon the result obtained using only the MW satellite population by E. O. Nadler et al. (2025), $m_{\mathrm{FDM}}>1.4\times 10^{-20}~\mathrm{eV}$, by about 24%. The precision of our limit is also comparable to that from the Lyman-$\alpha$ forest analysis in K. K. Rogers & H. V. Peiris (2021), which yields $m_{\mathrm{FDM}}>2.0\times 10^{-20}~\mathrm{eV}$.
@@
Besides, to obtain a more prior-independent constraint, we also adopt a posterior ratio of 20:1, corresponding to the point where the posterior probability drops to one-twentieth of its maximum value (D. M. Powell et al., 2023; R. E. Keeley et al., 2024; C. Y. Tan et al., 2025a). Our lower prior bound on $M_{\mathrm{hm}}$, $10^{7}~M_{\odot}$, coincides with the region where the posterior distribution flattens out, and we have verified the robustness of the 20:1 posterior ratio result by varying the prior range. For the FDM case, this criterion yields $M_{\mathrm{hm}}<4.17\times 10^{7}~M_{\odot}$, corresponding to
$m_{\mathrm{FDM}}>1.42\times 10^{-20}~\mathrm{eV}\ (\text{20:1}\ \mathrm{posterior\ ratio}).$ (19)
@@
Our result is about three times stronger than the posterior odds ratio of 20:1 constraint derived from strong gravitational lensing by D. M. Powell et al. (2023), who obtained $m_{\mathrm{FDM}}>4.4\times 10^{-21}~\mathrm{eV}$. We note that D. M. Powell et al. (2023) and C. Y. Tan et al. (2025a) adopt a posterior odds ratio, while E. O. Nadler et al. (2021a) use a likelihood ratio. However, because our posterior flattens toward the lower prior boundary and we adopt a uniform prior in $\log_{10}(M_{\mathrm{hm}})$, these criteria reduce to the same effective threshold as our posterior ratio.
@@
We present the full PDFs and contour maps of the MW and M31 parameters for the FDM case in Appendix A. We have verified that, aside from the shared parameters, the posterior distributions of the other parameters are consistent with those obtained from the separate MW and M31 fits. Furthermore, we rerun our model using the larger M31 mass prior from E. Patel & K. S. Mandel (2023), and obtain similar and slightly weaker constraint results with $m_{\mathrm{FDM}}>1.68\times 10^{-20}\,\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.34\times 10^{-20}\,\mathrm{eV}$ (20:1 posterior ratio).
V.2 Warm dark matter